Purifying Mg2+ from sea water. Aqueous sodium hydroxide is added to an aqueous solution that contains 0.0099 M magnesium chloride and 0.021 M calcium chloride. What will be the concentration of the metal ion that precipitates first (as M(OH)2) at the time the second metal ion begins to precipitate? Ksp (Mg(OH)2) = 6.3 x 10-10 and Ksp (Ca(OH)2) = 6.5 x 10-6
Hint: Use Ksp values to determine the solubility of each product to determine the order of precipitation.
Answer #1

Given : concentration of Mg2+ = [Mg2+] = 0.0099 M

concentration of Ca2+ = [Ca2+] = 0.021 M

(a.) The dissociation equation of Mg(OH)2 is : Mg(OH)2 (s) Mg2+ (aq) + 2 OH- (aq)

Let molar solubility of Mg(OH)2 = s

[Mg2+] = [Mg(OH)2] + concentration of Mg2+

[Mg2+] = s + 0.0099 M

[OH-] = 2 * [Mg(OH)2]

[OH-] = 2s

Ksp [Mg(OH)2] = [Mg2+][OH-]2

6.3 x 10-10 = (s + 0.0099 M) * (2s)2

Assuming s << 0.0099 M

6.3 x 10-10 = (0.0099 M) * (2s)2

6.3 x 10-10 = (0.0099 M) * (4s2)

s2 = (6.3 x 10-10) / (0.0099 M)

s2 = 6.36 x 10-8

s = (6.36 x 10-8)1/2

s = 2.52 x 10-4 M

[OH-] = 2s

[OH-] = 2 * (2.52 x 10-4 M)

[OH-] = 5.04 x 10-4 M

(b.)

The dissociation equation of Ca(OH)2 is : Ca(OH)2 (s) Ca2+ (aq) + 2 OH- (aq)

Let molar solubility of Ca(OH)2 = s

[Ca2+] = [Ca(OH)2] + concentration of Ca2+

[Ca2+] = s + 0.021 M

[OH-] = 2 * [Ca(OH)2]

[OH-] = 2s

Ksp [Ca(OH)2] = [Ca2+][OH-]2

6.5 x 10-6 = (s + 0.021 M) * (2s)2

Solving for x, s = 7.545 x 10-3 M

[OH-] = 2s

[OH-] = 2 * (7.545 x 10-3 M)

[OH-] = 0.0151 M

Since less concentration of OH- is required to precipitate Mg(OH)2, therefore, Mg(OH)2 will precipitate first and Ca(OH)2 will precipitate second.

concentration of OH- when Ca2+ begins to precipitate = 0.0151 M

Ksp Mg(OH)2 = [Mg2+][OH-]2

6.3 x 10-10 = [Mg2+] * (0.0151 M)2

[Mg2+] = (6.3 x 10-10) / (0.0151 M)2

[Mg2+] = 2.8 x 10-6 M